Find p-adic norm of |6|_3


1
Expert's answer
2021-05-26T14:08:26-0400

The "p-"adic norm of a rational number "\frac{a}{b}=p^\alpha\frac{a'}{b'}" where "p\nmid a', b'" is given by "p^{-\alpha}". As "6=3\times 2" and "3" does not divide "2", we have "|6|_3=3^{-1}=1\/3" .